KeigoTakami:

you guys wanna help me one more time- Solve x3 = 64 over 27. (4 points) A-±8 over 3 B-8 over 3 C-±4 over 3 D-4 over 3 You guys are good at math how-

kayla233:

UH

kayla233:

im gonna try to figure it out

kayla233:

idk how but I will

dude:

\(x^3=\dfrac{64}{27}\) Do the cube root of each number (numerator and denominator) Do you know how to find the cube root?

KeigoTakami:

mmhm

dude:

Cool Just do \(\sqrt[3]{64}\) and \( \sqrt[3]{27}\) for now